OpenStudy (melstutes):

A cell phone tower in a field casts a shadow of 100 ft. At the same time, a 4 ft, 6 in. post near the tower casts a shadow of 3 ft, 4 in. Find the height of the tower in feet.

OpenStudy (anonymous):

You have 100 ft shadow from the tower so we have x/100 and then we have 4.6 post with a shadow of 3.4 so we have 4.6 / 3.4 Now we transform the .4 from 3.4 into feet and .6 in 4.6 into feet .4 is 4/12 = 1/3 = .333333 of a foot. .6 is 6/12 = 1/2 = .5 so now we have 4.5 / 3.33 Now set up the problem as a proportion \( \huge \frac{x}{100} = \frac{4.5}{3.33} \) Do you got it from here?

OpenStudy (melstutes):

135?????

OpenStudy (anonymous):

Yes correct!!!!

OpenStudy (melstutes):

Thank you very much